# R3-RAG: Learning Step-by-Step Reasoning and Retrieval for LLMs via Reinforcement Learning

### v1

@@ 6 Conclusion
We propose R3-RAG, a framework that leverages reinforcement learning to teach LLMs how to reason and retrieve step by step, optimizing the interleaved reasoning and retrieval process with both outcome rewards and fine-grained process rewards through relevance-based document verification. Experimental results demonstrate that R3-RAG significantly outperforms extensive iterative RAG methods across multiple benchmarks. Further analyses show the effectiveness of R3-RAG’s reward design and its transferability across different retrievers.
@@
Despite the effectiveness of R3-RAG, our approach has two main limitations. First, our experiments rely on publicly available academic datasets, which may lack sufficient diversity and realism, potentially limiting the model’s robustness and generalization to real-world user queries. Future research could address this by curating more diverse and representative datasets. Second, our cold-start data quality is constrained by the reasoning capabilities of the foundation model (DeepSeek-v3). Due to its limitations, we could not reliably generate correct reasoning trajectories for certain challenging multi-hop questions, even after repeated sampling. Using stronger models, such as DeepSeek-R1, could further enhance cold-start data quality in future work.
@@
Our research complies with ethical standards in NLP research. We utilize publicly available datasets (HotpotQA, 2WikiMultiHopQA, and MuSiQue) and models (Llama-3.1-8B, Qwen2.5-7B) with appropriate licenses for academic research. Our derivative models maintain compatibility with their original licenses. Detailed information about datasets, models, frameworks, and licensing can be found in Appendix H.
@@
Regarding data privacy and content safety, we conducted a thorough review of the datasets to ensure they do not contain personally identifying information or offensive content. The datasets consist of questions and answers derived from Wikipedia, which undergoes content moderation. The questions focus on factual knowledge and do not target sensitive personal information. No additional data collection involving human subjects was conducted for this research, eliminating risks related to personal data exposure. Furthermore, we did not implement any mechanisms that could potentially generate harmful or offensive content. The training objectives were focused on improving factual accuracy and reasoning capabilities for answering multi-hop questions based on retrieved information.
Appendix A Data Generation Details of Cold Start
The complete prompt templates used in Algorithm 1 are shown in Figure 6 and Figure 7. The process consists of two main parts. In the first step as illustrated in Figure 6, the model analyzes the problem comprehensively, planning the solution trajectory and deciding whether to generate a retrieval query or provide a final answer based on information sufficiency. For subsequent reasoning steps presented in Figure 7, the prompt incorporates previous reasoning, queries, and retrieved documents, guiding the model to evaluate retrieved information, identify knowledge gaps, and determine whether additional retrieval is needed or if sufficient information has been gathered to answer the question.
After obtaining an initial dataset through Algorithm 1, to further enhance data quality, we vary the model’s sampling temperature to produce multiple samples for each problem. We then evaluate these different solution paths using Best-of-N (BON) selection, ultimately choosing the highest quality trajectory that correctly arrives at the correct answer, adheres to the proper format, and uses the most concise reasoning steps possible. Through this process, we obtain the dataset $\mathcal{D}_{CS}$ finally used for cold start.
Algorithm 2 Training Data Generation
0:  Trajectory-augmented dataset $\mathcal{D}_{CS}$
0:  Imitation learning training dataset $\mathcal{D}_{CS\_train}$
1:  $\mathcal{D}_{CS\_train}\leftarrow\emptyset$
2:  for each $(q_{i},a_{i}^{*},T_{i})\in\mathcal{D}_{CS}$ do
3:     $\mathtt{context}\leftarrow q_{i}$
4:     $\mathtt{history}\leftarrow\emptyset$
5:     for $j=1$ to $|T_{i}|$ do
6:        $\mathtt{input}\leftarrow\mathtt{context}\cup\mathtt{history}$
7:        if $j<|T_{i}|$ then
8:           $(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})\leftarrow s_{i}^{j}$
9:           $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,\psi_{i}^{j},\,d_{i}^{j})$
10:        else
11:           $(\xi_{i}^{j},\,a_{i})\leftarrow s_{i}^{j}$
12:           $\mathtt{output}\leftarrow\mathtt{Format}(\xi_{i}^{j},\,a_{i})$
13:        end if
14:        $\mathtt{history}\leftarrow\mathtt{history}\cup\{\mathtt{output}\}$
15:        $\mathtt{pair}\leftarrow(\mathtt{input},\mathtt{output})$
16:        $\mathcal{D}_{CS\_train}\leftarrow\mathcal{D}_{CS\_train}\cup\{\mathtt{pair}\}$
17:     end for
18:  end for
19:  return  $\mathcal{D}_{CS\_train}$
For each trajectory in $\mathcal{D}_{CS}$, we generate multiple training examples where each example consists of the original question concatenated with all previous reasoning steps as input, and the current reasoning step as output. This approach naturally addresses the masking requirement of retrieval documents, as retrieved documents are included in the input context during training, and the input context is masked during the training process. We define this processed training dataset as $\mathcal{D}_{CS\_train}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m}$, where $m=\sum_{i=1}^{|\mathcal{Q}^{\prime}_{CS}|}|T_{i}|$. Algorithm 2 illustrates our approach to converting the trajectory-augmented dataset $\mathcal{D}_{CS}$ into this training dataset $\mathcal{D}_{CS\_train}$. For each data $(q_{i},a_{i}^{*},T_{i})$ in $\mathcal{D}_{CS}$, we format the structured data $s_{i}^{j}$ in each step of $T_{i}$ into strings that conform to our required format. For a trajectory $T_{i}$, we iterate through each step, using all steps preceding the current one as the input for the training dataset, and the content of the current step as the output, ultimately creating $|T_{i}|$ training examples.
Appendix B Training Details of Cold Start
Based on the training dataset $\mathcal{D}_{CS\_train}$ generated from Algorithm 2, we implement Supervised Fine-Tuning (SFT) to train our model. Applying the maximum likelihood estimation principle, the loss function can be formulated as
$\mathcal{L}_{CS}=-\sum_{(x,y)\in\mathcal{D}_{CS\_train}}\log P(y|x;\theta),$ (6)
where $x$ is equal to $\mathtt{input}$ representing the combination of the question and all previous reasoning steps of the question, $y$ is equal to $\mathtt{output}$ representing the current step output of the question, and $\theta$ denotes the model parameters. When expanded across all questions in our trajectory-augmented dataset $\mathcal{D}_{CS}$, this loss function becomes
$\mathcal{L}_{IR}=-\sum_{i=1}^{|\mathcal{Q}^{\prime}_{CS}|}\sum_{j=1}^{|T_{i}|}\log P(s_{i}^{j}|h_{i}^{j};\theta),$ (7)
where $h_{i}^{j}=(q_{i},s_{i}^{1},\ldots,s_{i}^{j-1})$ represents the history of steps preceding step $j$ for question $i$, with the convention that $s_{i}^{0}=\emptyset$ for all $i$. The trained model is called as $\mathcal{R}^{3}$-RAG-CS.
Appendix C Online Data Sampling of RL
@@
We use multi-hop question-answering datasets $\mathcal{Q}_{RL}$ that are distinct from the cold-start datasets ($\mathcal{Q}_{RL}\cap\mathcal{Q}_{IL}=\emptyset$) for reinforcement learning, with online sampling logic consistent with the cold-start phase. We denote the sampled dataset as $\mathcal{D}_{RL}=\{(q_{i},a_{i}^{*},T_{i})\mid(q_{i},a_{i}^{*})\}$ and the final training dataset as $\mathcal{D}_{RL\_train}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m^{\prime}}$, where $m^{\prime}=\sum_{i=1}^{|\mathcal{Q}_{RL}|}|T_{i}|$, with each pair consisting of the question with previous reasoning steps as input and the current reasoning step as output.
@@
Appendix D Generalized Advantage Estimation and KL Divergence Definition
@@
Appendix E Implementation Details
@@
Appendix F Details of Analysis for Retriever Quality
HotpotQA 2WikiMultiHopQA MuSiQue Avg
Methods Retriever ACC F1 EM ACC F1 EM ACC F1 EM ACC
Llama-3.1-8B
@@ Appendix F Details of Analysis for Retriever Quality
Appendix G Potential Risk
@@
Appendix H Details of Research Artifacts and Licenses
@@ H.1 Datasets and Licenses
  • HotpotQA Yang et al. (2018) - Distributed under the CC BY-SA 4.0 license
@@
  • MuSiQue Trivedi et al. (2022) - Distributed under the CC BY 4.0 license
@@
H.2 Models, Frameworks, and Their Licenses
@@
The retrieval query: [If information is insufficient, generate one query based on the analysis.]
@@
The final answer: [If information is sufficient, provide the final answer based on the analysis.]
@@
#### Current Input:
The question: {question}
@@
# Role
You are an expert in LLMs and RAG, with extensive expertise in decomposition.
@@
The problem analysis: [Analyze information based on previous steps and question]
@@
- **Utilizing parameter knowledge**: If the retrieval results of the query are not relevant, check if the model’s own parameter knowledge can response the query.
@@
#### Current Input:
The question: {question}. The Previous Steps:{steps_formatted.strip()}

### v2

@@ 6 Conclusion
We propose R3-RAG, a framework that leverages reinforcement learning to teach LLMs how to reason and retrieve step-by-step, optimizing the interleaved reasoning and retrieval process with both outcome rewards and fine-grained process rewards through relevance-based document verification. Experimental results demonstrate that R3-RAG significantly outperforms existing iterative RAG methods across multiple benchmarks. Further analyses show the effectiveness of R3-RAG’s reward design and its transferability across different retrievers.
@@
Despite the effectiveness of R3-RAG, our approach has two main limitations. First, our experiments rely on publicly available academic datasets, which may lack sufficient diversity and realism, potentially limiting the model’s robustness and generalization to real-world user queries. Future research could address this by curating more diverse and representative datasets. Second, our cold-start data quality is constrained by the reasoning capabilities of the foundation model (DeepSeek-v3). Due to the model’s limitations, we could not reliably generate correct reasoning trajectories for certain challenging multi-hop questions, even after repeated sampling. Using stronger models, such as DeepSeek-R1, could further enhance cold-start data quality in future work.
@@
Our research complies with ethical standards in NLP research. We utilize publicly available datasets (HotpotQA, 2WikiMultiHopQA, and MuSiQue) and models (Llama-3.1-8B, Qwen2.5-7B) under appropriate licenses for academic research. Our derivative models comply with their original licenses. Detailed information about datasets, models, frameworks, and licensing can be found in Appendix F.
@@
Regarding data privacy and content safety, we conducted a thorough review of the datasets to ensure they do not contain personally identifiable information (PII) or offensive content. The datasets consist of questions and answers derived from Wikipedia, which undergoes content moderation. The questions focus on factual knowledge and do not target sensitive personal information. No additional data collection involving human subjects was conducted for this research, eliminating risks related to personal data exposure. Furthermore, we did not implement mechanisms that could generate harmful or offensive content. The training objectives were focused on improving factual accuracy and reasoning capabilities for answering multi-hop questions based on retrieved information.
Acknowledgements
We would like to thank the anonymous reviewers for their insightful comments. This work was supported by the Shanghai Pilot Program for Basic Research - Fudan University 21TQ1400100 (22TQ018).
Appendix A Online Data Sampling of RL
@@
We use multi-hop question-answering datasets $\mathcal{Q}_{RL}$ that are distinct from the cold-start datasets ($\mathcal{Q}_{RL}\cap\mathcal{Q}_{IL}=\emptyset$) for reinforcement learning, with online sampling logic consistent with the cold-start phase. We denote the sampled dataset as $\mathcal{D}_{RL}=\{(q_{i},a_{i}^{*},T_{i})\mid(q_{i},a_{i}^{*})\}$ and the final training dataset as $\mathcal{D}_{\mathrm{RL}}^{\mathrm{train}}=\{(\mathtt{input}_{k},\mathtt{output}_{k})\}_{k=1}^{m^{\prime}}$, where $m^{\prime}=\sum_{i=1}^{|\mathcal{Q}_{RL}|}|T_{i}|$, with each pair consisting of the question with previous reasoning steps as input and the current reasoning step as output.
@@
Appendix B Generalized Advantage Estimation and KL Divergence Definition
@@
Appendix C Implementation Details
@@
Appendix D Details of Analysis for Retriever Quality
[…]
@@ Appendix D Details of Analysis for Retriever Quality
Appendix E Potential Risk
@@
Appendix F Details of Research Artifacts and Licenses
@@ F.1 Datasets and Licenses
  • HotpotQA Yang et al. (2018) - Distributed under the $C$ BY-SA 4.0 license
@@
  • MuSiQue Trivedi et al. (2022) - Distributed under the $C$ BY 4.0 license
@@
F.2 Models, Frameworks, and Their Licenses
@@
The retrieval query: [If information is insufficient, generate one query.]
@@
The final answer: [If information is sufficient, provide the final answer.]
@@
#### Current Input: The question: {question}
@@
# Role: You are an expert in large language models and knowledge retrieval.
@@
The problem analysis: [Analyze information based on previous steps and the question]
@@
- **Utilizing parameter knowledge**: If the retrieval results of the query are not relevant, check if your parameter knowledge can response the query.
@@
#### Current Input: The question: {question}. The Previous Steps:{steps_formatted}